OpenStudy (anonymous):

What is the magnitude of the energy of a bond formed between a potassium (K+) cation and an iodide (I−) anion? The ionic radii of K+ and I−, are 152 pm and 206 pm, respectively. Assume the Born exponent n is 10. Please report your answer in joules.

OpenStudy (aaronq):

use the borne-lande equation

OpenStudy (anonymous):

i used but i can not found answer

OpenStudy (anonymous):

who knows answer

OpenStudy (anonymous):

in borne-lande equation what the value of distance between ions

OpenStudy (aaronq):

add up the radii

OpenStudy (anonymous):

i used this formula BE= (Z*e^2/4*pi*epsilon_0*r0)*(1-1/n) but i can not find answer.

OpenStudy (anonymous):

can you solve it

OpenStudy (anonymous):

Magnesia (MgO) is a very stable ionic solid commonly used as an inert containment vessel in high temperature experiments. The Born exponent, n, is 5.6 for MgO and its Madelung constant is 1.75. The ionic radii of Mg2+ and O2− are 48 and 145 pm, respectively. (a) Calculate the dissociation energy of a single MgO ion pair in the gaseous state. Express your answer in units of electron volts (eV) per ion pair.(b) Calculate the cohesive energy of an MgO crystal. Express your answer in units of kJ/mol. can you solve this please

OpenStudy (aaronq):

post what you did and i'll check it

OpenStudy (anonymous):

E=(1.602*10^-19)^2*+1*(-1)*6.023*10^23/4*31.4*8.85*10^-12*358*10^-12*(1-1/10) IN EQUATION WHERE IS THE WRONG STATEMENT

OpenStudy (aaronq):

\(E=\dfrac{(1.602*10^{-19})^2*(6.023*10^{23})M}{4\pi*(8.85*10^{-12})*(358*10^{-12})}(1-\dfrac{1}{10}) \) this reduces to: \(E=349362.16 *M~J/mol~\) but you need to know the madelung constant, M, i looked around but couldnt find it.

OpenStudy (anonymous):

thanks

OpenStudy (aaronq):

no problem

OpenStudy (anonymous):

What is the value of M in this sisuation

OpenStudy (anonymous):

Anybody got the answer?